# Demystifying Platform Requirements for Diverse LLM Inference Use Cases

### v1

@@ V-B Parallelization Strategies
Figure 8: Comparing different parallelism strategies for LLM inference on HGX:H00x8.
We show runtime breakdown in Fig. 9 to understand these performance gains. We ran GPT3-175B with batch size 8, beam size 4, and 4000 input tokens and 256 output tokens. In the compute-bound prefill phase, the compute of all NPUs is fully saturated, and thus, GEMM and SA times remain constant, irrespective of micro-batch size. The differentiating factor is the collective times. TP:8 requires two ARs for each layer, thus a total of 192 collectives with 8 NPUs. Meanwhile, PP:8 all needs to send messages at the interface of pipeline stages, thus only 7 collectives. In the memory-bound decode phase, we see a contrasting behavior. Due to the very small AR message size, the collective time remains small. Having smaller micro-batch sizes with PP further reduces collective time. With PP, serialized execution of layers results in larger latency. It should be noted that while total latency would be higher, per-stage latency would be reduced as PP goes down.
Figure 9: GPT3 runtime breakdown for various parallelism strategies on HGX: H100x8.
VI Platform Characteristics Analysis with GenZ
In this section, we use GenZ to study different platform characteristics. We explore four key platform characteristics: TFLOPS ($HW_{c}$), Memory BW ($BW_{fast}$), Interconnect link bandwidth ($BW_{link}$), Interconnect link latency ($T_{Link}$) To understand the effect of each characteristic on LLM inference performance, we vary each parameter keeping everything else constant. We assume an NPU with 800 TFLOPS, 40 GB HBM memory @ 4000 GB/s memory bandwidth, 300 GB/s interconnect bandwidth, and link latency of 2 us. For each parameter, we study all use cases in Table IV and all models in Table I, but in the interest of space we only show a few key data points in the main text.
(a) Prefill
(b) Decode
Figure 10: Scaling TFLOPS for various models on Rag-based question answering use case. Mixtral requires only 2 accelerators. GPT3 uses 12 accelerators with TP:12.
VI-A TFLOPS Scaling
We vary the platform TFLOPS by 2$\times$, 4$\times$, 8$\times$, 12$\times$. We show Mixtral 8x7B and GPT-3 inference on platforms with 2 and 12 NPUs, respectively. We chose RAG-based question answering as the representative use case of the study. Fig. 10 shows the runtime breakdown for the prefill and decode stage. We observe 4.33$\times$ and 2$\times$ reduction in prefill latency going 12$\times$ the original compute TFLOPS. For GPT3-175B, we see a smaller improvement in TTFT latency as it is dominated by the communication time, but the increased FLOPS would only reduce the prefill and attention time.
$\bullet$ Increasing computation has the potential to reduce the prefill latency significantly, but there is no impact on the decoding latency. $\bullet$ The benefits from increasing FLOPS diminish as platform size grows due to the significant contribution of network latency in prefill.
VI-B Memory Bandwidth Scaling
We vary the platform memory bandwidth by 2$\times$, 4$\times$, 8$\times$, 12$\times$. We show Mixtral 8x7B and GPT-3 inference on the platform with 2 and 12 NPUs, respectively. We chose text summarization as the representative use case for the study. Fig. 11 shows the runtime breakdown for the prefill and decode stage. We observe 4.21x and 1.73x reduction in decode latency when going to 8x the original compute memory BW. Similar to TFLOPS scaling, we see a smaller improvement in decode latency when it is dominated by the communication time.
$\bullet$ Increasing bandwidth has the potential to reduce the decoding latency significantly, but there is no impact on the prefill latency. $\bullet$ Like in TFLOPS, the benefits from increasing fast memory bandwidth diminish as platform size grows, owing to the substantial impact of network latency during decoding.
(a) Prefill
(b) Decode
Figure 11: Scaling Memory bandwidth per accelerator for text summarization use case. Mixtral requires only 2 accelerators. GPT3 uses 12 accelerators with TP:12.
VI-C ICN BW
Next, we vary the platform interconnect network bandwidth by 1$\times$, 2$\times$, 3$\times$, 6$\times$. We show GPT3-175B and GPT4-175B inference on the platform with 24 and 84 NPUs, respectively. We chose code generation as the representative use case of this study. Fig. 12 shows the runtime breakdown for the prefill and decode stages. We observe 1.44x and 1.56x reduction in prefill latency as ICN BW is increased to 6x the original BW.
$\bullet$ Increasing ICN BW has the potential to reduce the prefill latency significantly, but there is minimal impact on the decoding latency. $\bullet$ The benefits from increasing interconnect network bandwidth are only visible when platform size is significant such that network latency is in prefill.
(a) Prefill
(b) Decode
Figure 12: Scaling interconnect network BW for platform running code-generation use case. GPT3 and GPT4 uses 24(TP:12+PP:2) and 84(TP:21+PP:4) NPUs respectively.
VI-D ICN Link Latency
Finally, to study the impact of platform interconnect network link latency, we reduce the $T_{link}$ remove 2 us to 0.1 us. GPT3-175B and GPT4-175B inference on the platform with 24 and 84 NPUs, respectively. We choose code generation as the representative use case of this study. Fig. 13 shows the runtime breakdown for the prefill and decode stage. We observe a 1.8x and 2.9x reduction in the decoding latency as $T_{link}$ goes from 2 us to 0.1 us.
Our interconnect network modeling for the platform assumes the optimal topology for linking accelerators. Delving into different topologies for LLM inference presents a vast design space and currently a topic of research [127].
$\bullet$ Reducing interconnect network latency reduces the decode latency significantly, but there minimal impact on the prefill latency. $\bullet$ The benefits from reducing interconnect network link latency are only visible when platform sizes are of a significant size such that network latency is in decode.
(a) Prefill
(b) Decode
Figure 13: Scaling Interconnect network link latency for a model with code-generation use case. GPT3 uses 24 accelerators with TP:12+PP:2, and GPT4 uses 84 accelerators with TP:21+PP:4.
VI-E Fewer Chips with offload.
The most effective approach for serving inference involves storing all model weights and key-value (KV) cache in the fast memory of the NPU, such as HBM. However, this may not always be feasible due to resource limitations. In such cases, the model weights and KV cache can be transferred to slower memory. For LLM inference, where both model weights and KV cache are utilized simultaneously, a viable strategy is to offload a fixed portion of model weights and KV cache. Illustrated in Fig. 14 shows profiling of LLaMA3-70B with parameters $S_{b}=4,\tau_{p}=8000,\tau_{d}=256$ on platforms equipped with 2 and 4 A100-40GB GPUs (PCIe) and consequences of employing an offload-based system for LLM inference are depicted. While the model fits on the platform at smaller batch sizes, an increase in batch size leads to greater KV cache requirements, resulting in the offloading of some parameters. Consequently, inference latency rises by 13.2x and 12.7x due to offloading for 2xA100 and 4xA100 configurations, respectively. Thus offloading parameters can ensure correctness, but comes with a heavy penalty in the inference performance.
Figure 14: Running LLaMA3-70B with some parameters offloaded to slower memory via the host. All points on the right side of the red line of the plot have offloaded parameters.
VII Platform Requirements Estimation with GenZ
@@
VII-A Platform Memory Capacity Requirement
@@
As model sizes increase, the ratio of KV cache parameters to active weights diminishes progressively. Specifically, the ratio of the largest KV cache size (Code Gen) to active weights is 82%, 11%, 20%, 27%, and 2.8% for LLaMA2-7B, Mixtral-8x7B, LLaMA2-70B, GPT3-175B, and GPT4-1.8T, respectively. Moreover, MoE models exhibit significantly smaller KV caches compared to dense LLMs. Additionally, it is noteworthy that the KV cache expands linearly with the batch size. Furthermore, Mixtral 8x7B and LLaMA3-70VB employ GQA to minimize the KV cache footprint, thus enhancing batch size capabilities.
@@
VII-B Platform Compute Requirement
Given that the prefill stage is predominantly compute-bound, the platform prerequisites for TFLOPs are determined by prefill stage across all scenarios. Fig. 16(a) shows the platform-level compute TFLOPs required to meet the requirements in Table IV for various models.
@@ VII-B Platform Compute Requirement
VII-C Platform Memory Bandwidth Requirement
@@
VIII Extreme Scale Requirements: AI assistant

### v2

@@ IV-C Optimal Parallelization Strategies for MoEs
Figure 12: Comparing different parallelism strategies for Mixtral-8x22B model inference on HGX:H00x8. $S_{b}=4,\tau_{p}=4k,\tau_{d}=256,Chunk=512$.
V Impact of Model Architectures on HW Scaling
LLM architecture plays a pivotal role in determining their computational efficiency, scalability, and inference behavior. Model architectures are constantly evolving to optimize for specific challenges, such as memory usage, computational cost, and performance across diverse tasks. We compare the four most prominent LLM architectures, covering all SOTA open-source released models.
Traditional Transformers (Dense): Original foundational models like Transformer [44], GPT, and BERT have a fully connected attention mechanism where every token attends to every other token. This architecture have quadratically scaling compute and memory with input sequence length.
Dense Transformers with Group-query Attention: Group-query attention (GQA) [79] mitigates rising memory costs by sharing key-value caches across heads. This optimizes attention sparsity while maintaining accuracy, improving efficiency for long-context tasks.
Mixture of Experts (MoE): MoEs leverage sparsity by activating only a subset of ”expert” sub-networks for each token.
Mamba: The Mamba architecture combines attention mechanisms with flavor of RNNs. Mamba focuses on optimizing memory and compute utilization, making it particularly advantageous for large-scale deployments.
To understand the impact of these model architectures in the various stages of LLM inference, we compare a model of each type of comparable size. We pick LLama2-7B  [80], LLama-3.1-8B [52], Mixtral-8x7B (12.8 Active)  [46], and Falcon-Mamba-7b  [81] for our comparison. We study the impact of increasing context length & batch size on the latency of stages.
1
Impact on prefill stage: As the context grows (LABEL:fig:model_arch_impact(a)), All models exhibit linear scaling, for mamba model even though cache size is constant, they still need to have initial scan step linearly. With custom implementation of scan kernels, growth could be made sub-linear  [58]. For growing batch sizes (LABEL:fig:model_arch_impact(b)), all architectures show a linear increase in latency. Dense, Dense-GQA, and MoE scales maintain more efficient scaling. We don’t see any particular benefit in the Dense-GQA since the model is mainly compute-bounded, and having fewer KV heads doesn’t have any effect on the runtime.
2
Impact on decode stage: For increasing context lengths (LABEL:fig:model_arch_impact(c)), Mamba’s runtime doesn’t change since the generation time is context length independent. Dense sees a significant rise as context length increases due to the quadratic cost of dense attention. The effect on dense-GQA and MoE is mitigated since the KV cache growth factor is much smaller. Growing batch size (LABEL:fig:model_arch_impact(d)), Dense-GQA, mamba, and MoE grows almost linearly, which shows batching can help in increasing the throughput. For the dense model, the growth is linear, but the slope is higher due to the larger size of the KV cache.
3
Impact on chunked stage: In chunked prefill, a chunk of 512 tokens is constructed with multiple decode batches. Rest of the tokens are taken as a small chunk from the prefill request. Increasing context lengths for the decode batches,(LABEL:fig:model_arch_impact (e)), similar to the decode stage, larger context length equates to larger KV caches, and thus latency increases linearly. For batch sizes scaling (LABEL:fig:model_arch_impact (f)), Dense-GQA and MoE have a low growth rate due to fewer KV heads leading to smaller cache size. An interesting observation is that the MoE model has a larger latency compared to the dense model since the prefill portion of the chunk will ensure all experts are activated and thus would lead to higher latency.
VI Platform Requirements Estimation
@@
VI-A Platform Memory Capacity Requirement
@@
As model sizes increase, the ratio of KV cache parameters to active weights diminishes progressively. Specifically, the ratio of the largest KV cache size (Code Gen) to active weights is 82%, 11%, 20%, 27%, and 2.8% for LLaMA2-7B, Mixtral-8x7B, LLaMA2-70B, GPT3-175B, and GPT4-1.8T, respectively. Moreover, MoE models exhibit significantly smaller KV caches compared to dense LLMs. Additionally, it is noteworthy that the KV cache expands linearly with the batch size. Furthermore, Mixtral 8x7B and LLaMA3-70B employ GQA to minimize the KV cache footprint, thus enhancing batch size capabilities.
@@
VI-B Platform Compute Requirement
Given that the prefill stage is predominantly compute-bound, the platform prerequisites for TFLOPs are determined by prefill stage across all scenarios. Fig. 19(a) shows the platform-level compute TFLOPs required to meet the requirements in Table III for various models.
@@ VI-B Platform Compute Requirement
VI-C Platform Memory Bandwidth Requirement
@@
VII Studies on Future AI Inference Platform Design
In this section, we demonstrate the usefulness of GenZ in generating insights for future AI inference platform design.
(a) TFLOPS Scaling
(b) Memory Bandwidth Scaling
(c) Interconnect Network Bandwidth Scaling
(d) Interconnect Network Link Latency Scaling
Figure 20: Impact of scaling individual HW characteristics on runtime of different stages. The red points in the plot represent the optimal points for the stage and workload. The fixed NPU parameters are FLOPs=2PFLOPs, Memory=360 GB @ 12 TB/s and ICN link with 500 us link latency and 1.8 TB/s bandwidth.
VII-A Case Study I: Scaling HW Characteristics
We study how improving different HW characteristics impacts the latency of different stages. For all our experiments, we assume a hypothetical Dense-GQA-5T parameter model defined in Table IV. We explore four key hardware characteristics: TFLOPS, Memory BW , ICN link bandwidth, ICN link latency. To understand the effect of each characteristic on LLM inference performance, we vary each parameter in an isolated fashion. We vary the context length $\in\{1k,32k\}$, and for running with chunking; we vary the chunk size $\in\{384,1536\}$ and batch size $\in\{1,256\}$. Table VI summarizes this case study’s findings and Fig. 20 quantifies the observed trends.
VII-A1 TFLOPS Scaling
We vary the platform TFLOPS to increase the compute-to-memory BW ratio(C:M ratio). A higher C:M ratio benefits the operators with large arithmetic intensity due to additional compute units. The prefill stage, especially with long context, gets a very good reduction in latency as C:M increases. For smaller context, the improvement stops after C:M reaches 2000. The decode doesn’t improve as expected due to all layers being memory-bound. For chunking, the larger chunk sizes with fewer decode batches has the highest improvement with an increase in the C:M ratio. With more decode batches, or the smaller chunk size, there is only limited gain from the TFLOPs increase.
VII-A2 Memory Bandwidth Scaling
Next, we increase the memory bw of each NPU in the AI platform, thus reducing the C:M ratio. The prefill stage has no benefit from memory BW boost since it is completely compute-bound. The decode stage latency drops proportionally to memory BW increase since the decode stage is traditionally memory-bounded. For the chunking, the primary benefit is for scenarios where decode batches have accumulated and may constitute a significant portion of the overall runtime. For all other cases, there is no improvement in the chunk processing latency due to memory bandwidth improvement.
VII-A3 ICN Bandwidth Scaling
Our studies were done with a 32-NPU AI platform. Thus, communication latency composes a significant portion of the total latency. To further reduce the communication latency component, we scale the ICN BW. We see that the prefill stage benefits tremendously from bandwidth scale-up since the large message sizes in the prefill are primarily ICN bandwidth bound. Decode generally has much smaller message sizes, O($\approx$ 50-100 kBs); thus, bandwidth scale-up doesn’t improve its performance. With chunking, the message sizes are proportional to the chunk sizes. Thus, cases with large chunk-size benefit the most from the ICN BW increase. With chunk size, the benefit is limited since the collective might not be as significant a portion as other layers.
VII-A4 ICN Link Latency Scaling
Finally, we study the impact of platform interconnect network link latency; we reduce the $T_{link}$ remove 2.5 us to 0.1 us. For prefill, the communication component in a large context is smaller, and thus, reducing link latency does not significantly impact the run-time but is helpful for smaller context workloads. In the decoding process, link latency constitutes the majority of the total time involved. Therefore, a significant reduction in link latency leads to a notable decrease in decode latency. For chunking, reducing link latency has a more pronounced effect on smaller chunk sizes, while its impact on larger chunk sizes is minimal.
Characteristic Prefill Stage Decode Stage Chunked-Prefill
TFLOPS
$\uparrow$ Large (long
context)
$\uparrow$ Small (short
context)
(Memory- bound)
$\uparrow$ Large (large chunks,
fewer batches)
$\uparrow$ Small (small chunks,
more batches)
Memory
BW
(Compute- bound) $\uparrow$ Proportional
$\uparrow$ Only when decode
accumulates
ICN BW $\uparrow$ Large (ICN- bound large messages) (Small messages: 50-100kBs)
$\uparrow$ Large chunks
$\uparrow$ Small (Small chunks)
ICN Link Latency
$\uparrow$ Small (short
context)
(large context)
$\uparrow$ Significant
$\uparrow$ Small chunks
$\uparrow$ Small (Large chunks)
TABLE VI: Improvement in LLM Inference at different stages as we scale different platform characteristics.
Configuration
Platform
Inspiration
Compute/node
PFLOPS
Memory per node
(SRAM/HBM)
Network
Scale-up Scale-out
Peak Platform
Power(kW)
Multiple
GPUs
Nvidia-GB200 [82] 4.5
128MB@40 TB/s
192GB@8 TB/s
Switch{8}@900GB/s
Switch{4}@900GB/s
57.2
Single
SRAM wafer
Cerebras-CS3 [83] 125
44GB@21 PB/s
12TB@14.6 TB/s
On-Wafer@214 PB/s
23
Multiple
SRAM chips
Groq-GroqChip [84] 0.75
256MB@80 TB/s
FC{64}@3.2TB/s
Ring{16}@256GB/s
276.8
Multiple
ASICs
Etched-Soho [39] 45
256MB@80 TB/s
192GB@8 TB/s
Switch{8}@900GB/s
Switch{4}@900GB/s
96
TABLE VII: Platform architecture designs for comparison. Each of our platform’s architectures is inspired by a current system architecture. Power of 4 x DGX B200 platform(8xB200 GPU) [82]. Assumed power of a larger ASIC platform (NPUs have 10x flops of B200 GPU). Power of 16 GroqRack[85]. The architectures in the table are mapped to GenZ parameters defined in Section III-B, Section III-C. Column 3 is $FLOPS$, col 4 has BWmem, Capmem, BWsram, Capsram, col 5 has scaling dimension along with BWlink for each network level.
Figure 21: Comparing normalized throughput and throughput/energy of different platform architectures across various workloads. ‘X’ represents that the architecture is out of memory when running that workload. */* represents the platform architecture with the best performance. We run all workloads with batch size 4, $\tau_{d}=1024$, and chunk size 512.
VII-B Case Study II: Comparing Diverse Platforms
LLM inference throughput is heavily correlated to the characteristics of the AI platform used to serve LLMs. How to build next-generation LLMs is an open research question. Even commercially, industrial giants have hedged their bets on different platform architectures. We can categorize these architectural choices into four key buckets:
(1) Multiple GPUs: Traditional general-purpose SIMD or dataflow machines connected with a memory cache on-chip and connected to large memory banks. Examples of these are GPUs [86, 87, 88, 89, 90], TPUs [91, 92], and other AI accelerators  [93, 94, 95, 96].
(2) Wafer-scale chips with uber-fast on-wafer interconnect connecting cores and having very large on-chip SRAM, connected to large off-chip memory  [83].
(3) Multiple SRAM chips: Cluster of small chipset-based accelerators with large on-chip SRAM without any back-up memory.  [84, 97].
(4) Transformer-specific ASICs with a very large number of compute cores with a small memory cache on-chip and connected to large memory banks [39].
Using GenZ , we compare four representative platform architectures defined in Table VII running various workloads for current models (8B, 70B, 405B) and future model architectures (5T, 10T). Since the size and architecture of these platforms are completely different, there are different costs associated with each system. As we can’t estimate the cost of all components, we use energy consumed as a proxy and Tokens/kWH as the comparison metric. We calculate the energy used in running workloads on different platforms. The energy consumption of each platform is modeled as a linear function of the utilization of its individual components  [98]. We consider four main power components: Static/Idle, Compute, Memory, and Network. For each operator, the total energy is:
$E_{\text{op}}=T_{\text{op}}\times\left(P_{\text{static}}+P_{\text{C}}\cdot U_{\text{C}}+P_{\text{mem}}\cdot U_{\text{mem}}+P_{\text{icn}}\cdot U_{\text{icn}}\right)$ (2)
where $E_{\text{op}}$/$T_{\text{op}}$ are the operator energy/execution time, $P_{\text{static}}$ is the static power, and $P_{\text{C}},P_{\text{mem}},P_{\text{icn}}$ represent the peak power consumption of the compute, memory and network components, respectively. The terms $U_{\text{C}},U_{\text{mem}},U_{\text{icn}}$ denote their corresponding utilization factors for given operator. We use $P_{\text{static}}:P_{\text{C}}:P_{\text{mem}}:P_{\text{icn}}::3:4:2:1$  [^5] Future studies can use more fine-grained energy modeling simulators such as Accelergy [99], AccelWatch [98]. .
Fig. 21 compares the normalized throughput and normalized throughput/energy of four platform architectures across various workloads and stages of LLM inference. We run all workloads with batch size 4 and $\tau_{d}=1024$. For the chunked stage, we use a chunk size of 512. For LLama3 & GPT-3, we use TP=8 for GPUs and ASIC, TP=64, and PP=16 for SRAM chipsets. The use case indicates the model that was running and the input context size. For GPT-4, we use TP=32 for GPUs and ASIC. A summary of their performance is as follows: 1 GPUs: Excel in most decode and chunked workloads when the model cannot fit into SRAM, benefiting from high aggregate memory bandwidth. 2 Single SRAM Wafer: Leads in prefill workloads across most use cases due to superior energy efficiency. Performs best in decode and chunked stages when the entire model and weights fit within SRAM. 3 SRAM Chiplets: Optimal when model fits on chips but are consistently outperformed by other architectures in perf/energy, primarily due to their high power consumption owing to their very large platform size O(100s chips). 4 Transformer ASICs: Thrive in high compute demands, especially for future larger models/large context lengths.
$\bullet$ Choosing platform architecture for Performance: SRAM wafer/SRAM chips provide the best performance when the model fits on SRAM due to superior on-chip memory BW. ASIC is best in the prefill stage for very large models. In decode and most chunked workloads, ASIC and GPU give similar performance due to similar memory bandwidth.
$\bullet$ Choosing platform architecture for Perf/energy: SRAM Wafer is best for prefill stage and decode/chunked when the entire model fits on SRAM as lower energy used for running a single chip compared to multiple racks of NPUs, GPUs are best for running the rest of the decode and chunked workloads due to lower energy consumption accored to denser ASIC chip. If the ASIC chip can run with similar or lower power than GPU, it would be the best choice for achieveing highest performance/kWH in future larger models.
VII-C Case Study III: Exploring HBD Design Choices
As model sizes continue to scale, the minimum number of NPUs required to meet stringent Service Level Objectives (SLOs) has grown steadily. For instance, running GPT4, demands 64 H100 GPUs. Looking ahead, this baseline is poised to rise further, necessitating careful planning not just for NPU hardware but also for the accompanying network architecture—a challenge in its own right.
The concept of a high-bandwidth domain (HBD), defined as a group of NPUs interconnected via high-bandwidth scale-up links, plays a pivotal role here. Nvidia has progressively expanded the size of HBDs, scaling from 8 NPUs in the DGX H100 system to an impressive 72 NPUs in the GB200 NVL72. This case study delves into determining the optimal HBD size and designing effective interconnections between HBDs. To this end, we explore and compare various network architectures across a 256-NPU setup. We use the same individual NPU with 9 PFlops compute and 256 GB HBM, providing 13.5 TB/s for all configurations. Table VIII shows the bandwidth and link latency of different interconnect types that we consider for building the platforms. We keep the network topology fixed as 2 levels of switch followed by a ring for the third dimension. Using these, we build five different network architectures shown in Table IX. We ran models with TP=64 and PP=4.
Fig. 22 shows the throughput of different network configurations across running different workloads on 256 NPUs. Config D with all NPUs connected to a single HBD has the highest throughput but would also be the most costly to build. In contrast, Config B, with 64 NPUs per HBD, gives a similar throughput for the prefill stage at a much lower cost. Config E, with 64 NPUs per HBD connected via optical links as scale-out interconnect, achieves throughput that is comparable to config D at a lower cost for all stages.
High BW (SL) InfiniBand (IB) Optical
Latency (ns) 500 10000 200
BW (GB/s) 1800 256 900
TABLE VIII: Link latency and bandwidth for different interconnects. High bandwidth(SL) is a scale-up link similar to NVlink, UAlink, etc.  [100, 101, 102, 90, 103, 104, 105, 106, 107].
Config NPU Counts ICN Type HBD Size
A 8, 8, 4 SL, IB, IB 8
B 8, 8, 4 SL, SL, IB 64
C 8, 16, 2 SL, SL, IB 128
D 8, 8, 4 SL, SL, SL 256
E 8, 8, 4 SL, SL, Optical 64
TABLE IX: Comparison of different network configurations
Figure 22: Comparing throughput for different network config.
VII-D Case Study IV: Exploring Different NPU Microarchitectures and Offloading Choices
We compare systems with identical platform architectures but differing NPU microarchitectures, using SCALE-sim. The NPUs employ a systolic array with weight-stationary dataflow and spatial mapping. Keeping the total MAC units constant, we evaluate three configurations: (A) A single 256×256 systolic core. (B) Four 128×128 cores. (C) Four 128×128 cores with CPU offloading for MHA (Logit + Softmax + Attend) and KV cache storage. CPU has 8 TOPS and GPU to CPU link of 128 GB/s using PCIe. All NPUs are connected to a single 16GB HBM3e stack operating at 1.2 TB/s. Fig. 23 compares the prefill latency of these systems running LLaMA 3-8B (BF16) across varying input context sizes. System B achieves the lowest latency due to finer-grained kernel scheduling. System C, despite performance degradation from CPU offloading as the KV cache grows, can handle longer sequences, unlike Systems A and B, which are limited by fixed memory size. This analysis highlights GenZ’s capability to model diverse microarchitectures and offload compute.
Figure 23: Comparing different micro-architectures choices and compute offloading running LLama-3-8B prefill stage.
VII-E Extreme Scale Requirements: AI assistant